In a survey of 100 out-patients who reported at a hospital one day, it was found that 70 complained of fever,50 has stomach ache and 30 were injured. All 100 patients had at least one of the complaints and 44 had exactly two of the complaints. How many patients had all the three complaints?

Answer by ikleyn(52779)

.
Let F denotes the set of those who complain fever (and the number of elements in this set, at the same time).

Let S denotes the set of those who complain stomach ache (and the number of elements in this set, at the same time).

Let I denotes the set of those who complain injuries (and the number of elements in this set, at the same time).


Let FS denotes the intersection of the sets F and S (and the number of elements in this set, at the same time).

Let FI denotes the intersection of the sets F and I (and the number of elements in this set, at the same time).

Let SI denotes the intersection of the sets S and I (and the number of elements in this set, at the same time).


Finally, let FSI be the intersetion of the three sets F, S, and I (and the number of elements in this set, at the same time).


Then the number of those who has exactly 2 complaints is

    44 = FS + FI + SI - 3*FSI.     (1)


We have also the second equation from the condition


    100 = F + S + I - (FS + FI + SI) + FSI.     (2)


Now substitute the given values F= 70, S= 50  and  I = 30 into equation (2)


    100 = 70 + 50 + 30 - (FS + FI + SI) + FSI


and simplify it to get

    50 = (FS + FI + SI) - FSI.     (3)


Last step is to subtract eq(1) from eq(3). You will get


     6 = 2*FSI,


which implies  FSI = 6/2 = 3.


Answer.  The number of patients who have all the three complaints is 3.

Solved.

---------------

Regarding the formula (2), see the lesson
    - Advanced problems on counting elements in sub-sets of a given finite set
in this site.